# Analytical bounds for decoy-state quantum key distribution with discrete phase randomization

### v1

@@ VI Summary
In this work, we presented an analytical framework for evaluating the performance of QKD protocols based on discrete-phase randomization (DPR). Existing DPR-based security proofs often rely on computationally intensive numerical optimization [cao2015discrete, cao2020discrete], which limits their practicality in real-time or resource-constrained environments. We derived analytical bounds for the secret key rates of BB84 and MDI QKD protocols under DPR. These bounds closely matched numerical results and significantly reduced processing time, particularly when the number of phase slices was sufficiently large. The proposed methods substantially reduce computational overhead and are applicable to other MDI-type protocols.
Future work may extend this framework to finite-key analysis. Preliminary studies on DPR BB84 and twin-field (TF) QKD [jin2024finite, wang2023finite] use semidefinite programming to obtain numerical bounds. Our analytical approach could make such analysis more tractable. Additionally, the mode-pairing QKD protocol [zeng2022mode], a promising member of the MDI QKD family, merits further investigation due to its improved implementation feasibility compared to TF QKD [lucamarini2018overcoming].
@@ Appendix A Analytical BB84 Bounds
In this appendix, we derive Eq. 23 and Eq. 25 in the main text. We use $\gamma$ to represent either $X$ or $Y$ basis.
First, we can rewrite $Q_{\gamma}^{\alpha}$ and $Q_{\gamma}^{\alpha}E_{\gamma}^{\alpha}$ as [ma2005practical]
@@
$+\sum_{k=1}^{D-1}(e^{\alpha_{1}}p_{k}^{\alpha_{1}}\epsilon_{\mu\alpha_{1}}+e^{\alpha_{2}}p_{k}^{\alpha_{2}}\epsilon_{\mu\alpha_{2}})$
@@
$+A_{\alpha_{1}\alpha_{2}}[e^{\mu}Q^{\mu}-e^{\mu}p_{0}^{\mu}Y_{\gamma,0}^{\mu}-e^{\mu}p_{1}^{\mu}Y_{\gamma,1}^{\mu}]$
@@
where, in the first inequality, we use Eq. 13 repeatedly to write all yield terms either as a function of intensity $\mu$ or 0. Note that the left hand side of this inequality is a function of observable parameters. In the next two inequalities, we have written the right hand side as a function of $Y_{\gamma,1}^{\mu}$, the parameter we want to bound, as well as other observables and constants. Consequently, the lower bound of $Y_{\gamma,1}^{\mu}$ can be obtained as shown in Eq. 23.
For the parameters related to errors, the upper bounds are derived similarly. To derive $W_{y,1}^{\mu U}$, we can use the following
[…]
Consequently, we can get an upper bound on $W_{y,1}^{\mu}$ as given by Eq. 25.
@@ Appendix B Analytical MDI QKD Bounds
In this appendix, we derive Eq. 27 and Eq. 31 in the main text. We set $\alpha=\beta=\nu$ in analysis and use $\gamma$ to represent either $X$ or $Y$ basis.
First, we can rewrite $Q_{\gamma}^{\alpha\beta}$ and $Q_{\gamma}^{\alpha\beta}E_{\gamma}^{\alpha\beta}$ as [ma2005practical]
@@
$S_{2}=$ $e^{\alpha+\beta}(\sum_{k=0}^{N-1}p_{k}^{\alpha}p_{0}^{\beta}Y_{\gamma,k0}^{\alpha\beta}-p_{0}^{\alpha}p_{0}^{\beta}Y_{\gamma,00}^{\alpha\beta})$
@@
where $T_{1}$ and $\epsilon_{\bar{s}}$ are defined in the main text.
@@
where all introduced constants are defined in the main text.
@@
In terms of $e^{p,\mu\mu}_{x,11}$ in the key rate formula, we first bound $e^{b,\mu\mu}_{y,11}$ based on $W_{y,11}^{\mu\mu}$. Similar to what we did in Eq. 42, we replace all the $Q$ terms with $QE$ terms, change the less than or equal sign to the greater than or equal sign, and change the signs of $\epsilon$ terms from plus signs to minus signs. Consequently, a lower bound on $Q_{y}^{\alpha\beta}E_{y}^{\alpha\beta}$ is given by
$e^{\alpha+\beta}Q_{y}^{\alpha\beta}E_{y}^{\alpha\beta}$
$\geq$ $e^{\alpha+\beta}(p_{0}^{\alpha}Q_{y}^{0\beta}E_{y}^{0\beta}+p_{0}^{\beta}Q_{y}^{\alpha 0}E_{y}^{\alpha 0}-p_{0}^{\alpha}p_{0}^{\beta}Q_{y}^{00}E_{y}^{00}+p_{1}^{\alpha}p_{1}^{\beta}W_{y,11}^{\mu\mu}$
$-(p_{0}^{\alpha}+p_{0}^{\alpha}p_{0}^{\beta})\epsilon_{\alpha 0}-(p_{0}^{\beta}+p_{0}^{\alpha}p_{0}^{\beta})\epsilon_{\beta 0}$
@@
$\geq$ $T_{2}+e^{\alpha+\beta}p_{1}^{\alpha}p_{1}^{\beta}W_{y,11}^{\mu\mu}-\epsilon_{\bar{s}},$ (44)
using which the upper bound of $W_{y,11}^{\mu\mu}$ is obtained in Eq. 31.

### v2

@@ VI Summary
In this work, we presented an analytical framework for evaluating the performance of QKD protocols based on discrete-phase randomization (DPR). Existing DPR-based security proofs often rely on computationally intensive numerical optimization [2, 3], which limits their practicality in real-time or resource-constrained environments. We derived analytical bounds for the secret key rates of BB84 and MDI QKD protocols under DPR. These bounds closely matched numerical results and significantly reduced processing time, particularly when the number of phase slices was sufficiently large. The proposed methods substantially reduce computational overhead and are applicable to other MDI-type protocols.
There are several directions to pursue future work. First, we can extend this framework to include finite-key analysis. Preliminary studies on DPR BB84 [10] and twin-field QKD [16, 30] use semidefinite programming to obtain numerical bounds. Our analytical approach could make such analysis more tractable. Additionally, our techniques can be applied to the mode-pairing QKD protocol [37], which is another member of the MDI QKD family with improved scaling. Another valuable direction for future work is extending our DPR analytical analysis to fully-passive QKD architectures [33, 15]. Unlike DPR, where the phase is drawn from a known discrete set, fully-passive sources generate continuous, spontaneously-emitted phases whose probability distribution is not analytically characterized. This makes it difficult to construct the DPR decomposition and the associated closed-form bounds. Progress in this direction would require explicit models for the intrinsic phase distribution and for the conditional statistics with respect to post-selection. Establishing such models could enable computationally efficient analytical security bounds for fully-passive QKD systems.
@@ Appendix A Analytical BB84 Bounds
In this appendix, we derive Eq. 23 and Eq. 25 in the main text. We use $\gamma$ to represent either $X$ or $Z$ basis.
First, we can rewrite $Q_{\gamma}^{\alpha}$ and $Q_{\gamma}^{\alpha}E_{\gamma}^{\alpha}$ as [18]
@@
$+e^{\alpha_{1}}p_{0}^{\alpha_{1}}\epsilon_{\alpha_{1}0}+e^{\alpha_{2}}p_{0}^{\alpha_{2}}\epsilon_{\alpha_{2}0}+e^{\alpha_{1}}p_{1}^{\alpha_{1}}\epsilon_{\mu\alpha_{1}}+e^{\alpha_{2}}p_{1}^{\alpha_{2}}\epsilon_{\mu\alpha_{2}}$
@@
$+A_{\alpha_{1}\alpha_{2}}[e^{\mu}Q^{\mu}_{\gamma}-e^{\mu}p_{0}^{\mu}Y_{\gamma,0}^{\mu}-e^{\mu}p_{1}^{\mu}Y_{\gamma,1}^{\mu}]$
@@
where $p_{k}^{\alpha}$ is the probability distribution function defined in Eq. 5, $\epsilon$ terms are given by Eq. 13, and $A_{\alpha_{1}\alpha_{2}}$ is a constant defined in Eq. 24. In the first inequality, we use Eq. 13 repeatedly to write all yield terms either as a function of intensity $\mu$ or 0. Note that the left hand side of this inequality is a function of observable parameters. In the next two inequalities, we have written the right hand side as a function of $Y_{\gamma,1}^{\mu}$, the parameter we want to bound, as well as other observables and constants. Consequently, the lower bound of $Y_{\gamma,1}^{\mu}$ can be obtained as shown in Eq. 23.
For the parameters related to errors, the upper bounds are derived similarly. To derive $W_{x,1}^{\mu U}$, we can use the following bound:
$Q_{x}^{\alpha_{1}}E_{x}^{\alpha_{1}}$ $=p_{0}^{\alpha_{1}}W_{x,0}^{\alpha_{1}}+p_{1}^{\alpha_{1}}W_{x,1}^{\alpha_{1}}+\sum_{k=2}^{D-1}p_{k}^{\alpha_{1}}W_{x,k}^{\alpha_{1}}$
$\geq p_{0}^{\alpha_{1}}(W_{x,0}^{0}-\epsilon_{\alpha_{1}0})+p_{1}^{\alpha_{1}}(W_{x,1}^{\mu}-\epsilon_{\mu\alpha_{1})})$
$=p_{0}^{\alpha_{1}}Q_{x,0}^{0}E_{x,0}^{0}+p_{1}^{\alpha_{1}}W_{x,1}^{\mu}-p_{0}^{\alpha_{1}}\epsilon_{\alpha_{1}0}-p_{1}^{\alpha_{1}}\epsilon_{\mu\alpha_{1}}.$ (39)
Consequently, we can get an upper bound on $W_{x,1}^{\mu}$ as given by Eq. 25.
@@ Appendix B Analytical MDI QKD Bounds
In this appendix, we derive Eq. 27 and Eq. 31 in the main text. We set $\alpha=\beta=\nu$ in analysis and use $\gamma$ to represent either $Z$ or $X$ basis.
First, we can rewrite $Q_{\gamma}^{\alpha\beta}$ and $Q_{\gamma}^{\alpha\beta}E_{\gamma}^{\alpha\beta}$ as [18]
@@
$S_{2}=$ $e^{\alpha+\beta}(\sum_{k=0}^{D-1}p_{k}^{\alpha}p_{0}^{\beta}Y_{\gamma,k0}^{\alpha\beta}-p_{0}^{\alpha}p_{0}^{\beta}Y_{\gamma,00}^{\alpha\beta})$
@@
where $T_{1}$ and $\epsilon_{\bar{s}}$ are defined in Eq. 28 and Eq. 30 separately.
@@
where $\epsilon$ terms are given by Eqs. 19 and 30, and $A,B,C$ are constants defined in Eq. 29.
@@
In terms of $e^{p,\mu\mu}_{z,11}$ in the key rate formula, we first bound $e^{b,\mu\mu}_{x,11}$ based on $W_{x,11}^{\mu\mu}$. Similar to what we did in Eq. 44, we replace all the $Q$ terms with $QE$ terms, change the less than or equal sign to the greater than or equal sign, and change the signs of $\epsilon$ terms from plus signs to minus signs. Consequently, a lower bound on $Q_{x}^{\alpha\beta}E_{x}^{\alpha\beta}$ is given by
$e^{\alpha+\beta}Q_{x}^{\alpha\beta}E_{x}^{\alpha\beta}$
$\geq$ $e^{\alpha+\beta}(p_{0}^{\alpha}Q_{x}^{0\beta}E_{x}^{0\beta}+p_{0}^{\beta}Q_{x}^{\alpha 0}E_{x}^{\alpha 0}-p_{0}^{\alpha}p_{0}^{\beta}Q_{x}^{00}E_{x}^{00}$
$+p_{1}^{\alpha}p_{1}^{\beta}W_{x,11}^{\mu\mu}-(p_{0}^{\alpha}+p_{0}^{\alpha}p_{0}^{\beta})\epsilon_{\alpha 0}-(p_{0}^{\beta}+p_{0}^{\alpha}p_{0}^{\beta})\epsilon_{\beta 0}$
@@
$\geq$ $T_{2}+e^{\alpha+\beta}p_{1}^{\alpha}p_{1}^{\beta}W_{x,11}^{\mu\mu}-\epsilon_{\bar{s}},$ (46)
where $T_{2}$ and $\epsilon_{\bar{s}}$ are given in Eq. 32 and Eq. 30, respectively. The upper bound on $W_{x,11}^{\mu\mu}$ can then be calculated as in Eq. 31.